# On the Equivalence of Koopman Eigenfunctions and Commuting Symmetries

### v3

@@ IV Formula for Koopman Eigenfunctions on a Region of Attraction
where $\mathcal{D}=\left\{\mathbf{z}\in\mathbb{C}^{n}\mid z_{i}\in\frac{1}{\lambda_{i}}\log(\mathbf{w}_{i}^{*}\mathbf{x}),\,\mathbf{x}\in\mathbb{R}^{n}\right\}$. Denote by $\widetilde{\Phi}_{F}(t,\mathbf{z})$ the flow in $\mathbf{z}$ coordinates. As $\mathbf{z}\to\infty_{n}$, or equivalently as $t\to\infty$ along each trajectory, the dynamics asymptotically approach $\dot{\mathbf{z}}=1_{n}$, such that the commuting symmetries approach simple translations. In order to extend these symmetries to the region of attraction, recall the following lemma.
@@
Consider flow $\widetilde{\Phi}_{F}(t,\mathbf{z})$ and given $\mathbf{z}^{\prime}\in\mathcal{D},\mathbf{v}\in T_{\mathbf{z}^{\prime}}\mathcal{D}$. Define the vector field $\widetilde{E}(\mathbf{z})$ such that, at every point on $\widetilde{\Phi}_{F}(\mathbb{R},\mathbf{z}^{\prime})$,
@@
The proof is omitted.
Note that the flow asymptotically approaches the constant flow, $\dot{\mathbf{z}}=1_{n}$, and a set of commuting symmetries is given by the simple translations with the generators $\big{[}\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{n}\big{]}$. Their pushforwards are
@@
They commute with each other because they are pushforwards of the commuting local frame $\big{[}\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{n}\big{]}$ by the diffeomorphism $\lim_{t\to\infty}\widetilde{\Phi}_{F}(-t,\mathbf{z})$. Substituting (17) into (8), we obtain a set of conservative linearizing vector fields as
@@
The integrals of $\widetilde{X}_{i}(\mathbf{x})$ are given by
@@
where we added the constant t-t under the limit. By computing the exponential of the logarithms of Koopman eigenfunctions $\widetilde{m}_{i}(\mathbf{z})$ and changing back to $\mathbf{x}$ coordinates, we obtain the following formula for the principal Koopman eigenfunctions on the region of attraction $A$ [27, Defn. 4].
@@
with $F_{n}(\mathbf{x})=F(\mathbf{x})-\nabla F(\mathbf{x}_{0})\mathbf{x}$, which is the path-integral formula proposed by Deka et al. in [28, 29]. In comparison, the advantage of (20) is replacing the limit of an integral by the limit of a function. On the condition for convergence, compared to [28, 29] which prove convergence by assuming a small spectral gap, Assumption 1 requires that the linearized system is non-resonant, which holds for a more general class of systems, e.g., systems with time-scale separation. The non-resonance assumption here is necessary for the existence of the Poincaré-Dulac normal form around an equilibrium point [30, 27], and simple resonance example can be used to show that (20) generally fails to converge in the resonance case. However, it should be also noted that non-resonance is not necessary for the existence of smooth Koopman eigenfunctions on any open subset of the region of attraction bounded away from the equilibrium [25], where the flow-box theorem applies.
@@
Eq. (20) is a particular case of the computation formula in [31, Prop. 2] for the isostable, which is formally defined through an Laplace average. An isostable is a level set of a principal Koopman eigenfunction such that all points on the isostable converge to the equilibrium at the same exponential rate. While the isostable formula is well-known in practice, its convergence is usually not guaranteed. The symmetry characterization of it and the sufficient condition provided in Assumption 1 are both steps toward analytical characterization of Koopman eigenfunctions and isostables.
@@
To demonstrate the practicality of (20), we consider the example of the reverse-time van der Pol oscillator: Using MATLAB, we verify the limiting expression (20) by computing the left eigenvectors and evaluating it over a grid of initial conditions. Both numerical convergence and the eigenfunction property are verified. The magnitude and phase of the two complex-conjugate principal eigenfunctions are plotted on the set $1,1]^{2}$, as shown in Fig. 1.
Figure 1: Filled contour plots for the Koopman eigenfunctions of the reverse-time van der Pol oscillator with $\mu=0.5$ computed using (20). Trajectories are simulated from every point on a $0.004$-interval rectangular grid for 1010 s to achieve sufficient convergence ($1\%$ maximum relative difference compared to simulating for 88 s).
@@ V Conclusion
This paper establishes the equivalence between Koopman eigenfunctions and commuting symmetries as a special case of the equivalence between commuting and integrable local frames. This equivalence converts the problem of identifying Koopman eigenfunctions—traditionally approached by fitting integrable linearizing vector fields (e.g., EDMD)—into the problem of identifying commuting symmetry generators that form a local frame. We demonstrate the importance of this equivalence in analysis by deriving an explicit formula for the principal Koopman eigenfunctions on the region of attraction of a locally asymptotically stable equilibrium point. In doing so, we also demonstrate the correspondence between Koopman eigenfunctions and commuting symmetries on the region of attraction.
@@
Recall Cartan’s identity for $1$-forms:
@@
Let $X_{i}\mathrel{\mathop{\ordinarycolon}}=\nabla z_{i}$ so that $\theta^{i}(Y)=X_{i}^{*}(Y)$. Then
@@
(Conservative $\Rightarrow$ commuting). Conversely, if $X_{i}=\nabla z_{i}$ and $\theta^{i}\mathrel{\mathop{\ordinarycolon}}=dz_{i}$, let $\{E_{i}\}$ be the dual frame to $\{\theta^{i}\}$, so $\big{[}X_{1}\;\cdots\;X_{n}\big{]}^{*}=\big{[}E_{1}\;\cdots\;E_{n}\big{]}^{-1}$. Since $d\theta^{i}=0$,
@@
$\frac{d}{dt}\Big{(}e^{-\lambda_{i}t}g_{i}^{(N)}(\mathbf{y}(t))\Big{)}=$
$e^{-\lambda_{i}t}\Big{(}\nabla g_{i}^{(N)}(\mathbf{y}(t))^{*}\big{(}\mathbf{\Lambda}\mathbf{y}(t)+R^{(N+1)}(\mathbf{y}(t))\big{)}-\lambda_{i}g_{i}^{(N)}(\mathbf{y}(t))\Big{)}.$
@@
$\frac{d}{dt}\Big{(}e^{-\lambda_{i}t}g_{i}^{(N)}(\mathbf{y}(t))\Big{)}=e^{-\lambda_{i}t}\mathcal{O}(\|\mathbf{y}(t)\|^{\,N+1}).$
@@
By local asymptotic stability, there exists a compact $U_{0}\subset A$ contained in the normal-form chart, and $t_{K}>0$ such that $\Phi_{F}(t,\mathbf{x})\in U_{0}$ for all $\mathbf{x}\in K$, $t\geq t_{K}$. Writing $\tau=t-t_{K}$ and $\mathbf{x}^{\prime}=\Phi_{F}(t_{K},\mathbf{x})\in U_{0}$, the uniform convergence on $U_{0}$ and the finite $|e^{-\lambda_{i}t_{K}}|$ give the uniform convergence on $K$. Finally, to verify the eigenfunction property,

### v4

@@ IV Formula for Koopman Eigenfunctions on a Region of Attraction
We introduce the following assumption to ensure the existence of smooth Koopman eigenfunctions around $\mathbf{x}_{0}$.
@@
where $\mathcal{D}=\left\{\mathbf{z}\in\mathbb{C}^{n}\mid z_{i}\in\frac{1}{\lambda_{i}}\log(\mathbf{w}_{i}^{*}\mathbf{x}),\,\mathbf{x}\in\mathbb{R}^{n}\right\}$. Denote by $\widetilde{\Phi}_{F}(t,\mathbf{z})$ the flow in $\mathbf{z}$-coordinates. As $\mathbf{z}\to\infty_{n}$, or equivalently as $t\to\infty$ along each trajectory, the dynamics asymptotically approach $\dot{\mathbf{z}}=1_{n}$, such that the commuting symmetries approach simple translations. In order to extend these symmetries to the entire region of attraction, recall the following lemma.
@@
Consider the flow $\widetilde{\Phi}_{F}(t,\mathbf{z})$, and, given $\mathbf{z}^{\prime}\in\mathcal{D}$ and $\mathbf{v}\in T_{\mathbf{z}^{\prime}}\mathcal{D}$, define the vector field $\widetilde{E}(\mathbf{z})$ at every point in $\widetilde{\Phi}_{F}(\mathbb{R},\mathbf{z}^{\prime})$ such that
@@
The proof is omitted. In words, the pushforward $\widetilde{E}(\mathbf{z})$ of a tangent vector $\mathbf{v}$ at $\mathbf{z}^{\prime}$ along the trajectory of the system $\widetilde{\Phi}_{F}(t,\mathbf{z}^{\prime})$, defines a symmetry generator on $\widetilde{\Phi}_{F}(\mathbb{R},\mathbf{z}^{\prime})$.
Note that the flow at the limit $\mathbf{z}\to\infty_{n}$ asymptotically approaches the constant flow $\dot{\mathbf{z}}=1_{n}$, for which a set of commuting symmetries is given by the simple translations generated by $\big{[}\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{n}\big{]}$. Their pushforwards are
@@
By Lemmas 4 and 1, each $\widetilde{E}_{i}(\mathbf{z})$ commutes with the flow $\widetilde{\Phi}_{F}(t,\mathbf{z})$. Moreover, $\widetilde{E}_{i}(\mathbf{z}),\widetilde{E}_{k}(\mathbf{z})$ commute with each other because they are pushforwards of the commuting $\mathbf{e}_{i},\mathbf{e}_{k}$ by the diffeomorphism $\lim_{t\to\infty}\widetilde{\Phi}_{F}(-t,\widetilde{\Phi}_{F}(t,\mathbf{z}))$. Substituting (17) into (8), we obtain a set of conservative linearizing vector fields as
@@
The integrals of $\widetilde{X}_{i}(\mathbf{z})$ are given by
@@
where the constant t-t is added before $t$ is taken to the limit, allowing the convergence of the limit. Computing the exponential of the logarithms of Koopman eigenfunctions $\widetilde{m}_{i}(\mathbf{z})$ and changing back to $\mathbf{x}$-coordinates, we obtain the formula for the principal Koopman eigenfunctions [28, Defn. 4] on the region of attraction $A$ as stated below.
@@
with $F_{n}(\mathbf{x})=F(\mathbf{x})-\nabla F(\mathbf{x}_{0})\mathbf{x}$, which is the path-integral formula proposed by Deka et al. in [29]. In comparison, the advantage of (20) is replacing the limit of an integral by the limit of a function. On the condition for convergence, compared to [29] which proves convergence by assuming a small spectral gap, Assumption 1 requires that the linearized system is non-resonant, which holds for a broader class of systems, e.g., systems with time-scale separation. The non-resonance assumption here is necessary for the existence of the Poincaré-Dulac normal form around an equilibrium point [30, 28], and a simple resonance example can be used to show that (20) generally fails to converge in the resonance case. However, it should also be noted that non-resonance is not necessary for the existence of smooth Koopman eigenfunctions on any open subset of the region of attraction bounded away from the equilibrium [27], where the flow-box theorem applies.
@@
Eq. (20) is a particular case of the computation formula in [31, Prop. 2] for the isostable, which is formally defined through a Laplace average. An isostable is a level set of a principal Koopman eigenfunction such that all points on the isostable converge to the equilibrium at the same exponential rate. While the isostable formula is well-known in practice, its convergence is usually not guaranteed. The symmetry-based characterization and the sufficient condition provided in Assumption 1 are both steps toward analytical characterization of Koopman eigenfunctions and isostables.
@@
To demonstrate the practicality of (20), we consider the example of the reverse-time van der Pol oscillator. Using MATLAB, we verify the limiting expression (20) by computing the left eigenvectors and evaluating it over a grid of initial conditions. Both numerical convergence and the eigenfunction property are verified. The magnitude and phase of the two complex-conjugate principal eigenfunctions are plotted on the set $1,1]^{2}$, as shown in Fig. 1.
Figure 1: Filled contours of the two principal Koopman eigenfunctions for the reverse-time van der Pol oscillator with $\mu=0.5$, computed via (20). Trajectories are simulated with ode45 from a rectangular grid with $0.004$ spacing for 1010 s; this achieves $\leq 1\%$ maximum relative difference over the square domain compared with simulating for 88 s.
@@ V Conclusion
This paper establishes the equivalence between Koopman eigenfunctions and commuting symmetries as a special case of the equivalence between commuting and integrable local frames. This equivalence transforms the problem of identifying Koopman eigenfunctions—traditionally approached by fitting integrable linearizing vector fields, e.g., EDMD—into the problem of identifying commuting symmetry generators that form a local frame. We demonstrate the importance of this equivalence in analysis by deriving an explicit formula for the principal Koopman eigenfunctions on the region of attraction of a locally asymptotically stable equilibrium point. In doing so, we also demonstrate the correspondence between Koopman eigenfunctions and commuting symmetries on the region of attraction.
@@
Recall Cartan’s identity for one-forms:
@@
Let $X_{i}=\nabla z_{i}$ so that $\theta^{i}(Y)=X_{i}^{*}(Y)$. Then
@@
(Conservative $\Rightarrow$ commuting). Conversely, if $X_{i}=\nabla z_{i}$ and $\theta^{i}=dz_{i}$, let $\{E_{i}\}$ be the dual frame to $\{\theta^{i}\}$, so $\big{[}X_{1}\;\cdots\;X_{n}\big{]}^{*}=\big{[}E_{1}\;\cdots\;E_{n}\big{]}^{-1}$. Since $d\theta^{i}=0$,
@@
$\frac{d}{dt}\Big{[}e^{-\lambda_{i}t}g_{i}^{(N)}(\mathbf{y}(t))\Big{]}=$
$e^{-\lambda_{i}t}\Big{[}\nabla g_{i}^{(N)}(\mathbf{y}(t))^{*}\big{[}\mathbf{\Lambda}\mathbf{y}(t)+R^{(N+1)}(\mathbf{y}(t))\big{]}-\lambda_{i}g_{i}^{(N)}(\mathbf{y}(t))\Big{]}.$
@@
$\frac{d}{dt}\Big{[}e^{-\lambda_{i}t}g_{i}^{(N)}(\mathbf{y}(t))\Big{]}=e^{-\lambda_{i}t}\mathcal{O}(\|\mathbf{y}(t)\|^{\,N+1}).$
@@
By local asymptotic stability, there exists a compact $U_{0}\subset A$ contained in the normal-form chart, and $t_{K}>0$ such that $\Phi_{F}(t,\mathbf{x})\in U_{0}$ for all $\mathbf{x}\in K$, $t\geq t_{K}$. Writing $\tau=t-t_{K}$ and $\mathbf{x}^{\prime}=\Phi_{F}(t_{K},\mathbf{x})\in U_{0}$, the uniform convergence on $U_{0}$ and the finite $|e^{-\lambda_{i}t_{K}}|$ give the uniform convergence on $K$. Finally, to verify the eigenfunction property, it holds that